OpenStudy (anonymous):

Calculate the discriminant and use it to determine how many real-number roots the equation has. 3x2 – 6x + 4 = 0 A. three real-number roots B. no real-number roots C. one real-number root D. two real-number roots

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@JoKeR0331 @KevinM.Williams159 @Lady.Liv1776 @Legends @sleepyjess @k142 @dan815 @desiraem @Samus_Aran

OpenStudy (anonymous):

Three

OpenStudy (anonymous):

I think?

OpenStudy (anonymous):

can you tag someone else to see what they think to??

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

Can't be 3. the highest exponent is 2, so you there can only be 0, 1, or 2 solutions. The discriminant is b² - 4ac. - If b² - 4ac = 0, there is 1 solution. - If b² - 4ac is positive, there are 2 solutions. - If b² - 4ac is negative, there are 0 solutions.

OpenStudy (anonymous):

my answer whould be c?

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

No. What did you get for the discriminant.

OpenStudy (anonymous):

sorry, i meant to say d!:)

OpenStudy (loser66):

I am with peachpi. They asked you to find discriminant. You MUST have it first.

OpenStudy (anonymous):

i am not sure how to do that, will you helpme??

OpenStudy (anonymous):

For the equation ax² + bx + c = 0, the discriminant is b² - 4ac. For your equation, a = 3, b = -6, and c = 4.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So what's b²-4ac?

OpenStudy (anonymous):

um is that the abc equation?? @peachpi

OpenStudy (anonymous):

I'm not sure what the a,b,c equation is. I gave you the numbers above. You just need to plug them into the expression b²-4ac to get the discriminant.

OpenStudy (anonymous):

-6^2-4(3)(4) @peachpi

OpenStudy (anonymous):

is this correct? @peachpi

OpenStudy (anonymous):

yes and now you need to simplify it to a single number.

OpenStudy (anonymous):